Problem 7-18 Supernormal Growth [LO 1]

Biarritz Corp. is growing quickly. Dividends are expected to grow at a rate of 30 percent for the next three years, with the growth rate falling off to a constant 7 percent thereafter.

If the required return is 13 percent and the company just paid a dividend of $2.60, what is the current share price? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

I have answered the question below

Please up vote for the same and thanks!!!

Do reach out in the comments for any queries

Answer:

D1=(2.6*1.30)=3,38

D2=(3.38*1.30)=4.394

D3=(4.394*1.3)=5.7122

Value after year 3=(D3*Growth rate)/(Required rate-Growth rate)

=(5.7122*1.070)/(0.13-0.070)

= 101.8676 ~ 101.87